# MyFramelinecolor=black, outerlinewidth=.1pt, roundcorner=1pt, innertopmargin=nerbottommargin=nerrightmargin=5pt, innerleftmargin=-1pt, backgroundcolor=gray!10!white The First VoicePrivacy Attacker Challenge Evaluation Plan Version 1.0

### v1

@@ 1 Context
The first three VoicePrivacy Challenge editions [3, 4, 5, 6] focused on the development of voice anonymization systems. In particular, the systems submitted to the VoicePrivacy 2024 Challenge had to meet the following requirements: (a) output a speech waveform; (b) conceal the speaker identity at the utterance level; (c) not distort the linguistic and emotional content. The processed utterances sound as if they were uttered by another speaker, which we refer to as a pseudo-speaker. The pseudo-speaker is selected independently for every utterance, and it can be an artificial voice not corresponding to any real speaker. In practice, many voice anonymization systems select the pseudo-speaker or modify prosody in a random or semi-random way using a random number generator. A semi-informed attack model [7] was assumed, whereby attackers have access to the voice anonymization system (but not to the random numbers drawn by that system for each utterance, if any), and they seek to re-identify the original speaker behind each anonymized trial utterance. Specifically, an ECAPA-TDNN [8] automatic speaker verification (ASV) system was provided by the organizers and trained by the participants on data anonymized using their anonymization system. While this attack model is undeniably the most realistic to date, the provided attacker system is not its strongest possible implementation as it does not exploit spoken content similarities, specific pseudo-speaker selection strategies [9], or stronger ASV architectures [10], among others.
@@ 3 Data
The training set is the train-clean-360 subset of the LibriSpeech [11] corpus. Besides the provided anonymized training data, participants are allowed to use the original train-clean-360 data. In addition, participants are allowed to use other resources such as speech corpora and pretrained models to develop their attacker systems.
@@ 5 Baseline attacker system
As a baseline, we consider the attacker system used in the VoicePrivacy 2024 Challenge [4] (see Figure 1). The ASV system (denoted ) is an ECAPA-TDNN [8] with 512 channels in the convolution frame layers, implemented by adapting the SpeechBrain [24] VoxCeleb recipe to LibriSpeech, and it is trained on anonymized training data. For a given trial utterance and enrollment speaker, the attacker computes the average speaker embedding of all anonymized enrollment utterances from that speaker and compares it to the speaker embedding of the anonymized trial utterance. Results for this baseline attacker system are reported in Table 3444Note, that these EER results may vary for different runs of anonymization algorithms and ASV training due to randomness (i.e. for B3, for B5).. The code to train the baseline attacker systems for given anonymized data is available in the GitHub VoicePrivacy 2024 Challenge repository: https://github.com/Voice-Privacy-Challenge/Voice-Privacy-Challenge-2024.555See Step 2: Evaluation, run_evaluation.py
@@ 6 Evaluation rules
Participants can use the training resources and development datasets specified in Section 3 in order to train their system and tune hyperparameters. The use of additional speech data and models is allowed provided that they are publicly and freely available and declared by email to attacker.challenge@inria.fr before the data declaration deadline (see Table 4).
@@ 7 Registration and submission of results
Participants/teams are requested to register for the evaluation. Registration should be performed once only for each participating entity using the registration form. Participants will receive a confirmation email within 48 hours after successful registration, otherwise or in case of any questions they should contact the organizers: attacker.challenge@inria.fr.
@@
Each participant may submit scores/results for one or more attacker systems, each targeting all anonymization systems or only some of them. Each single submission should include the EER and corresponding ASV scores (for the development and evaluation data) obtained with the proposed attacker system for 4 trial lists (in the same format as generated by the baseline attacker system666Example: link):

### v2

# MyFramelinecolor=black, outerlinewidth=.1pt, roundcorner=1pt, innertopmargin=nerbottommargin=nerrightmargin=5pt, innerleftmargin=-1pt, backgroundcolor=gray!10!white The First VoicePrivacy Attacker Challenge Evaluation Plan Version 2.1

@@ 1 Context
The first three VoicePrivacy Challenge editions [3, 4, 5, 6, 7, 8] focused on the development of voice anonymization systems. In particular, the systems submitted to the VoicePrivacy 2024 Challenge had to meet the following requirements: (a) output a speech waveform; (b) conceal the speaker identity at the utterance level; (c) not distort the linguistic and emotional content. The processed utterances sound as if they were uttered by another speaker, which we refer to as a pseudo-speaker. The pseudo-speaker is selected independently for every utterance, and it can be an artificial voice not corresponding to any real speaker. In practice, many voice anonymization systems select the pseudo-speaker or modify prosody in a random or semi-random way using a random number generator. A semi-informed attack model [9] was assumed, whereby attackers have access to the voice anonymization system (but not to the random numbers drawn by that system for each utterance, if any), and they seek to re-identify the original speaker behind each anonymized trial utterance. Specifically, an ECAPA-TDNN [10] automatic speaker verification (ASV) system was provided by the organizers and trained by the participants on data anonymized using their anonymization system. While this attack model is undeniably the most realistic to date, the provided attacker system is not its strongest possible implementation as it does not exploit spoken content similarities, specific pseudo-speaker selection strategies [11], or stronger ASV architectures [12], among others.
@@ 3 Data
The training set is the train-clean-360 subset of the LibriSpeech [13] corpus. Besides the provided anonymized training data, participants are allowed to use the original train-clean-360 data. In addition, participants were allowed to propose other resources such as speech corpora and pretrained models before the deadline (13th October). Based on the suggestions received from the challenge participants, in this version of the evaluation plan, we publish the final list of training data and pretrained models allowed for training attacker systems. All the allowed resources are listed in Table LABEL:tab:data-models-final-list.
For some models, the provided link is a webpage listing multiple versions of the model. In this case, unless otherwise stated, all model versions available on that page before 15th October 2024 can be used by participants in training their attacker systems. Participants are allowed to use any existing software in the development and training of their attacker systems. If the software uses pretrained models, these models should be explicitly listed in this table or on the main page (readme) of the corresponding repository before 15th October 2024.
@@ 5 Baseline attacker system
As a baseline, we consider the attacker system used in the VoicePrivacy 2024 Challenge [4] (see Figure 1). The ASV system (denoted ) is an ECAPA-TDNN [10] with 512 channels in the convolution frame layers, implemented by adapting the SpeechBrain [48] VoxCeleb recipe to LibriSpeech, and it is trained on anonymized training data. For a given trial utterance and enrollment speaker, the attacker computes the average speaker embedding of all anonymized enrollment utterances from that speaker and compares it to the speaker embedding of the anonymized trial utterance. Results for this baseline attacker system are reported in Table 4. The code to train the baseline attacker systems for given anonymized data is available in the GitHub VoicePrivacy 2024 Challenge repository: https://github.com/Voice-Privacy-Challenge/Voice-Privacy-Challenge-2024.444See Step 2: Evaluation, run_evaluation.py
@@ 6 Evaluation rules
Participants can use the training resources and development datasets specified in Section 3 in order to train their system and tune hyperparameters.
@@ 7 Registration and submission of results
Participants/teams are requested to register for the evaluation. Registration should be performed once only for each participating entity using the registration form. Participants will receive a confirmation email within 48 hours after successful registration, otherwise or in case of any questions they should contact the organizers:
@@
Each participant may submit scores/results for one or more attacker systems, each targeting all anonymization systems or only some of them. Each single submission should include the EER and corresponding ASV scores (for the development and evaluation data) obtained with the proposed attacker system for 4 trial lists (in the same format as generated by the baseline attacker system555Example: link):